OpenStudy (anonymous):

3 = 9 - 1(24 - 2(9)) = 3(9) - 1(24). how does 3=9-1(24-2(9)) equal or become 3(9) -1(24)?

OpenStudy (anonymous):

Distributive property, combined with the commutative property. \[ 9 - 1(24 - 2(9)) =9-24+2(9)= 3(9) - 24=3\]The negative one distributes across the 24 and the product of -2 and nine. We gather like terms (terms with a common factor of nine) to get the three times nine term.